Please help me with this question:
The formula M=5k over 18 (fraction) changes speed of k km/h to a speed of M m/s. If k=72 what is the value of M?

Thank you very much.

Answer by Alan3354(69443)

The formula M=5k over 18 (fraction) changes speed of k km/h to a speed of M m/s. If k=72 what is the value of M?
-------
M = 5k/18 = 5*72/18
M = 20